# Shift Parallelism: Low-Latency, High-Throughput LLM Inference for Dynamic Workloads

### v1

@@ 4.3. Performance Benchmarks
Latency vs. throughput tradeoff
@@
Variations across context sizes.
@@
Shift Parallelism provides up to 2.45 faster generation than DP and a similar output latency to TP. Ideally, the output latency should not depend on the input context size in theory, but in practice TPOT increases with the input size (see Figure 10). The main reason is that each output token needs to read more number of tokens from the KV cache as input context grows, and eventually the system becomes memory bandwidth bound. TP and Shift Parallelism parallelizes the attention layer (Section 2.3), and hence the KV cache, providing memory bandwidth, mitigating the output latency for long inputs.
@@
Latency vs. Arrival Rate
To investigate the performance between extremely high and low traffic rates, we test Shift Parallelism across a wide range of intermediate traffic by varying the request arrival rates. We measure TTFT and TPOT of an individual request, which both increases with higher traffic, and calculate the completion time as . The question is, where does the tradeoff happens, and how well Shift Parallelism transitions from latency to throughput optimization?
@@ 4.4. Shift Parallelism in Production
We fully integrated Shift Parallelism in our existing production environment (see Section 3.4). Running efficiently in production is not only about parallelism, but it also requires a plethora of other state-of-the-art techniques. To that extent, we integrated Shift Parallelism with SwiftKV (Qiao et al., 2025) and speculative decoding (Wang et al., 2025; Oliaro et al., 2025) in our production environment.
Figure 12 shows that our combined production simultaneously achieves highest throughput (lowest cost) and lowest completion time—all in one deployment—outperforming the best open source systems optimized for each metric individually 222We enabled the best available speculative decoding for each framework. These experiments were run on data sets generated using real-world production traces to compute throughput, and a mixture of ShareGPT, HumanEval and SWEBench to measure latency. As a result, these results are representative of performance achievable in real-world deployments. For more details, see the evaluation methodology in the appendix..
4.5. Cost Breakdown
We analyze the cost of individual system components by taking away one component at a time. Figure 13 shows the resulting breakdown of time to process 1,920 requests with Llama-70B and Qwen-32B models on a single node. We clearly see that SP (and hence Shift Parallelism) has a lower communication cost than TP. On the other hand, we observe two unaddressed performance bottlenecks that are not related to Shift Parallelism:
Attention time grows significantly with the sequence size, and therefore reduces the combined throughput. Recent papers address this issue using sparse attention [cite] and it is out of scope of this paper.
@@ 4.6. Limitations and Future Work
To enable these models, we use the SP generalizations in Section 3: i) Llama-17B-16E barely fits into a single GPU and when SP=8 is used in the base config., there is no memory left in the KV cache to support large context sizes. To enable long contexts, we use Algorithm 2 for the base config. (SP=4, TP=2). ii), Qwen-30B-A3B suffers from scaling because it only has 4 KV heads, and we use Algorithm 1 to scale the model across SP=8 GPUs.
Figure 14 compares the throughput and latency performance in across the board. The models sorted from larger to smaller. The sparse models attains a higher throughput and lower latency than dense models, simply because they have less number of active parameters (see Table 4). Compared to TP, Shift parallelism shows excellent performance on both sparse models: Up to 50% higher throughput without increasing the latency.
The smallest model, Qwen-30B-A3B, attains a much higher throughput (250k tokens/sec.) than the other models with DP, and the throughput suddenly drops either with TP and SP, especially with smaller end of input sizes. The discrepancy is due to the vLLM’s parallelization overhead with small models that is discussed in Section 4.5.
The sparse models beg further investigation in the context of expert parallelism (EP). Specifically, there is no prior work that combines SP with EP to to further optimize sparse models, which we will leave as a future work.
@@ 6. Conclusion
LLM’s have diverse real-world applications that yields different traffic patterns that have different performance requirement. Fundamentally, existing parallelisms (TP and DP) for inference optimize either for latency (for interactive applications) or throughput (for batch workloads), but do not optimize for multiple applications in the same deployment.
For supporting dynamic workloads, we need to switch back-and-forth between parallelisms, yet cannot switch across TP and DP due to their KV cache mismatch. As a remedy, we bring in SP (Ulysses) that is originally applied to training for high throughput, and generalized it for inference, and providing the KV cache invariance across SP and TP by addressing corner cases. We call our solution Shift Parallelism, where a deployment has two configs. that switch back and forth across SP and TP.
Our extensive benchmarking show that Shift Parallelism addresses the latency vs. throughput tradeoff in a low cost way, providing up to 50% more throughput compared without losing latency across traffic rates. We fully, integrated Shift Parallelism along with other SoTA that we use in production, and open source our inference system (Rajbhandari et al., 2025).

### v2

@@ 4.3. Performance Benchmarks
4.3.1. Latency vs. throughput tradeoff
@@
4.3.2. Variations across context sizes.
@@
Shift Parallelism provides up to 2.45 faster generation than DP and a similar output latency to TP. Ideally, the output latency should not depend on the input context size in theory, but in practice TPOT increases with the input size (see Figure 13). The main reason is that each output token needs to read more number of tokens from the KV cache as input context grows, and eventually the system becomes memory bandwidth bound. TP and Shift Parallelism parallelize the attention layer (Section 2.3), and hence the KV cache, providing memory bandwidth, mitigating the output latency for long inputs.
@@
4.3.3. Latency vs. Arrival Rate
To investigate the performance between extremely high and low traffic rates, we test Shift Parallelism across a wide range of intermediate traffic by varying the request arrival rates. We measure TTFT and TPOT of an individual request, which both increases with higher traffic, and calculate the completion time as . The question is, where does the tradeoff happen, and how well Shift Parallelism transitions from latency to throughput optimization?
@@
4.4. Cost Breakdown
We analyze the cost of individual system components by taking away one component at a time. Figure 15 shows the resulting breakdown of time to process a batch of requests with Llama-70B and Qwen-32B models on a single node. We clearly see that SP (and hence Shift Parallelism) has a lower communication cost than TP. On the other hand, we observe two unaddressed performance bottlenecks that are not related to Shift Parallelism:
Attention time grows significantly with the sequence size, and therefore reduces the combined throughput. Recent papers address this issue using sparse attention (Child et al., 2019) and it is out of scope of this paper.
@@ 4.5. Shift Parallelism in Production
We fully integrated Shift Parallelism in our existing production environment (see Section 3.4). Running efficiently in production is not only about parallelism, but it also requires a plethora of other state-of-the-art techniques. To that extent, we integrated Shift Parallelism with SwiftKV (Qiao et al., 2025) and speculative decoding (Wang et al., 2025; Oliaro et al., 2025) in our production environment (Rajbhandari et al., 2025).
Figure 16 shows that our combined production simultaneously achieves highest throughput (lowest cost) and lowest completion time—all in one deployment—outperforming the best open source systems optimized for each metric individually 222We enabled the best available speculative decoding for each framework. These experiments were run on data sets generated using real-world production traces to compute throughput, and a mixture of ShareGPT, HumanEval and SWEBench to measure latency. As a result, these results are representative of performance achievable in real-world deployments.. The compounding effect of SwiftKV and speculative decoding are denoted.
@@ 4.6. Limitations and Future Work
To enable these models, we use the SP generalizations in Section 3: i) Llama-17B-16E barely fits into a single GPU and when is used in the base config., there is no memory left in the KV cache to support large context sizes. To enable long contexts, we use Algorithm 1 for the base config. . ii), Qwen-30B-A3B suffers from scaling because it only has 4 KV heads, and we use KV cache replication (Section 3.2) to scale the model across GPUs.
Figure 17 compares the throughput and latency performance in across the board. The models sorted from larger to smaller. The sparse models attain a higher throughput and lower latency than dense models, simply because they have fewer active parameters (see Table 4). Compared to TP, Shift parallelism shows excellent performance on both sparse models: Up to 50% higher throughput without increasing the latency.
The smallest model, Qwen-30B-A3B, attains a much higher throughput (225k tokens/sec.) than the other models with DP, and the throughput suddenly drops either with TP and SP, especially with smaller end of input sizes. The discrepancy is due to the vLLM’s parallelization overhead with small models that is discussed in Section 4.4.
The sparse models beg further investigation in the context of expert parallelism (EP). Specifically, there is no prior work that combines SP with EP to further optimize sparse models, which we will leave as a future work.
@@ 6. Conclusion
LLM’s have diverse real-world applications that yield different traffic patterns that have different performance requirements. Fundamentally, existing parallelisms (TP and DP) for inference optimize either for latency (for interactive applications) or throughput (for batch workloads), but do not optimize for multiple applications in the same deployment.
For supporting dynamic workloads, we need to switch back-and-forth between parallelisms, yet cannot switch across TP and DP due to their KV cache mismatch. As a remedy, we bring in SP (Ulysses) that is originally applied to training for high throughput, and generalized it for inference, and providing the KV cache invariance across SP and TP by addressing corner cases. We call our solution Shift Parallelism, where a deployment has two configurations. that switch back and forth across SP (base config.) and TP (shift config.).
Our extensive benchmarking shows that Shift Parallelism addresses the latency vs. throughput tradeoff in a low cost way, providing up to 50% more throughput compared without losing latency across traffic rates. We fully integrated Shift Parallelism along with other SoTA that we use in production, and open source our inference system.
Appendix A Artifact Appendix
A.1. Abstract
Our evaluation in Section 4 involves extensive results. For assessing the real-world benefit, we will focus on running open-source production traces (Section 4.2.2). The artifact compares vLLM’s built-in parallelisms (DP and TP) and the proposed parallelisms (SP and Shift Parallelism) using ArcticInference plug-in system that we also use in our production. For reproducibility, we included the instructions in the repository, and uploaded the cleaned traces into a permanent archive.
A.2. Artifact check-list (meta-information)
Algorithm: Multi-GPU parallelism strategies (DP, TP, SP, Shift Parallel) for LLM inference.
Program: vLLM v0.10.1 and the corresponding ArcticInference plug-in version (both are open source).
Model: We use two open-source models downloadable from Hugging Face.
Data set: Two open-source production traces. Clean traces for reproducibility are archived in DOI:
https://doi.org/10.5281/zenodo.18240909. -
Run-time environment: Linux, Python 3.10
Hardware: A node with 8H200 Nvidia GPUs w/ NVLinks.
Metrics: TTFT, TPOT, completion time
Output: vLLM’s benchmarking script prints the result.
Experiments: Running vLLM benchmarking scripts (server and client). We provide the scripts.
How much disk space required (approximately)?: 100 GB to store both models.
How much time is needed to prepare workflow (approximately)?: 15 mins
How much time is needed to complete experiments (approximately)?: 5 node-hours. Experiments can be easily done in parallel across two nodes.
Publicly available?:
https://github.com/snowflakedb/ArcticInference -
Code licenses?: Apache-2
A.3. Description
A.3.1. How to access
vLLM is open source (https://github.com/vllm-project/vllm).
A.3.2. Hardware dependencies
We used the hardware described in Section 4.1.1), yet a generic Nvidia DGX-H200 node is sufficient for 1:1 reproduction. Optimal configurations, and hence the results may look different another type of multi-GPU node, yet the conclusion should be the same.
A.3.3. Software dependencies
We depend on vLLM v0.10.1 and the corresponding ArcticInference plug-in (commit 5e08f0f).
A.3.4. Data sets
We use the following datasets (links given):
A.3.5. Models
The models are open-source in Hugging Face (links given):
- (1)
- (2)
These can be downloaded as huggingface-cli download <model> --local-dir <model>
A.4. Installation
We built our work on vLLM, therefore it is essential to make vLLM work first. The installation details can be found in https://github.com/vllm-project/vllm, yet it should be as easy as: pip install vllm==v0.10.1
A.5. Experiment workflow
The step-by-step instructions can be found in
ArcticInference/benchmark/reproducibility.First the output file of an experiment should be obtained. Then the plotting script should be used to plot TTFT, TPOT, and completion time. The scripts for running the experiments and plotting them as in Figure 9–10 are provided.
A.6. Evaluation and expected results
Shift parallelism is expected to have a lower TTFT, TPOT, and completion compared to TP and DP. The raw files can be plotted using plot.py. Additionally, the columns (TTFT, TPOT, Completion) can be sorted to obtain Figure 11.
![[Uncaptioned image]](x3.png)